3. If the circuit in Fig. 2 is built and operated, it will be found that the resistor becomes much hotter than the inductor, even though both components drop the exact same amount of voltage and carry the exact same amount of current:
339.4 Alternator 34 2 9.02 mH 480 VAC 60 Hz 3394 Fig. 2. Alternator driving a series loop with resistor and inductor. There is an identical voltage drop across each device, and they carry exactly the same current. Explain why there is such a remarkable difference in heat output between these two components, given their identical voltage drops and currents.
Answer #1

It is all based on concept of Active and Reactive Power

See Resistor is Power disspating Component and Consumes Active or Useful Power in Watts

Given By P= V*I =(IR)* I= I2 R watts  .

This Power is consumed and Disspated by the resistor continiously at same rate every instant of the voltage cycle .

While in the indcutor Power is circulated during each cycle from source to inductor ( mahnetic enrgy stored) and from inductor to source ( magnetic energy dissipated ) and this occurs very fast as frequency of this is 60 Hz or 60 cycles . So Power just circulates to and fro unlike resistor that have constant power disspation and hence resistor gets more heated and inductor remains less heated to due to circulating Or Reacttive Power present withing Inductor .